Displaying reports 221-240 of 1079.Go to page Start 8 9 10 11 12 13 14 15 16 End
Reports until 13:44, Wednesday 22 January 2025
Lab Infrastructure General
Daniel Grass - posted 13:44, Wednesday 22 January 2025 (12085)
Keys placed in B106

I put a command hook in B106 to hold various keys on key rings. I put my key to B150 and a key to the fire cabinet in B102B.

Vacuum Systems Pumps (GQuEST)
Daniel Grass - posted 17:09, Tuesday 21 January 2025 (12084)
Custom Plates to Attach POC001 Caster Wheels to Agilent Base Plate Made, Cleaned, and Attached

[Alex, Daniel]

I designed some custom plates that mount to the Agilent Base Plate and allow POC001 Caster Wheels to be attached. We had 4 of these caster wheels around from the SNSPD Dewer Cart. The purpose is to move the vacuum pump system along with the dewer.

I attached the wheels to the plate with the 4 M6 screws included with the wheels. I attached the plate to the Agilent Base Plate using a 3/8-24 hex bolt (and a lock washer) from the "inside" of the Agilent Base Plate. I also used a #10-32 set screw to help align the custom plate into the Agilent Base Plate since there is a rotational degree of freedom. 

See attached part file, STEP file, drawing for tapping, and photos.

Images attached to this report
Non-image files attached to this report
GQuEST Output Filter Cavities
Torrey Cullen - posted 15:16, Tuesday 21 January 2025 (12083)
Second AOM chassis problem solved

[Ian,Torrey]

We found why the second chassis wasn't working. In order to diagnose we first confirmed these new RF amplifiers were working by snipping the leads and powering it via a variable power source. A deflected beam was seen after AOM3. We tried a bunch of short testing and finally concluded that the the 5 pin power connector was rotated 90 degrees with respect to the original power cable (in that picture, original means the correctly functioning chassis, new means the newer one that was built that was not functioning correctly). This orientation doesn't have a preferred direction, but they do need to be the same. I believe we could fix this by quickly swapping the wires here and keeping this 90 rotation, but we decided to resolder this box to make it the same as the old one. This is better for the long term.

We plugged everything back in and now see a deflected beam on AOM3. The left side of the box is significantly under supplying the AOM (i.e. there is a deflected beam, but way less than the other side). We will diagnose this tomorrow. 

Images attached to this report
Equipment General
Alex Ramirez - posted 14:46, Friday 17 January 2025 (12081)
Vacuum Pump Systems

[Alex, Daniel]

We have moved the agilent vacuum pumps (3x) to B111A. The systems have been organized into stacks along the wall as seperate pump systems.

It was then decided that because the Dewar cart is inherently on wheels and capable of being mobile; wheels should be added to the pump system.

Thus, Daniel and I are now retrofitting the old casters from the original thorlabs cart to fit the base plate for the vacuum pump to enable it to move with the dewar when the time comes.

This should be ready sometime next week for pump down with the Dewar assembly.

Environmental Dust/Particulates
Daniel Grass - posted 13:04, Friday 17 January 2025 (12080)
Particle Counter Measurements January 17, 2025

A smoke smell has been reported earlier in the morning (althoguh I haven't smelled anything, potentially because I went downstairs at 11 AM), so I took some particle counts.

  B111A North small table B111A Outside clean rooms west of north small table B111B Control area B111B Outside clean room by northeast shelves B111B Inside clean room B111 Antechamber
0.3 um 0 242 197 142 1 2737
0.5 um 0 40 59 81 0 307
1.0 um 0 31 44 57

0

151

The 0.3 um counts outside the clean rooms are ~3x higher than Tuesday, January 14. The larger particle counts are also elevated.

Vacuum Systems General (GQuEST)
Daniel Grass - posted 19:25, Thursday 16 January 2025 - last comment - 18:05, Friday 17 January 2025(12079)
Flanges moved around and installed on Laser Filter Cavity and Demonstrator Interferometer Vacuum Cubes

[Jeff, Daniel]

We placed a 10" blank flange on the Input Laser Filter Cavity (LFC) Vacuum Cube North End and tightened the screws to 34 Nm. We moved 10" flange with electronics D-Subs to the Demonstrator Interferometer Vacuum Cube (transmitted arm) so that we can lock the interferometer with a piezo. I tightened the screws to 13.6 Nm with one pass but decided to go home and not work alone. These flanges and faces on the cubes were a bit dirty, so we did some cleaning with the wipes meant to clean inside vacuum chambers. We moved the 10" to 4.5" zero length reducer flange to the top of the Input LFC Vacuum Cube, but the screws did not want to go in more than ~0.35" before almost all of them had a lot of resistance. The holes seem tapped and deep enough; I measured 1.8" from the top of the flange to the bottom of the hole. I think we need to re-tap or clean up these threads. Ian recommended getting a thread chaser kit. I think this one would be a good option. We placed foil on the newly exposed face on the Input LFC Vacuum Cube. I want to seal up the flanges before potentially making chips on top of the vacuum chamber.

Comments related to this report
Daniel Grass - 18:05, Friday 17 January 2025 (12082)

I tightened the screws on the 10" electronics flange to 34 Nm on the IFO.

GQuEST SHG
Jeffrey Wack - posted 14:59, Wednesday 15 January 2025 (12078)
AOM troubleshooting

[Torrey, Jeff]

The second AOM was not working, and this was resolved by unplugging the BNC on the AOM and replugging it.

There was no 775 light reaching the second filter cavity, so we investigated the second AOM on the SHG sled. After confirming that it was receiving RF power correctly by swapping cables with the first AOM, we removed the cover and visually inspected the crystal to ensure it looks good. We added a pickoff mirror and a camera after the AOM and could see two spots and fiddling with alignment for a while we could still not get the power of the frequency shifted beam to an acceptable amount (we want to see roughly 50/50). Finally after unplugging and replugging the BNC the two spots were about equal.

We then had to realign the fiber coupler. We left it at

76 uW in, 40uW out -> 52% efficiency

Vacuum Systems General (GQuEST)
Daniel Grass - posted 16:37, Tuesday 14 January 2025 (12076)
2.75 in 1550 nm AR Coated Viewport installed on Laser Filter Cavity Output Vacuum Cube; Vacuum Cubes moved around

[Jeff, Daniel]

The clean room smelled less of smoke, so I took some more particle counts at 2:50 PM, after ~150 minutes of the air purifier being on in B111A, to ensure we can do vacuum work in B111B and to see how well the air purifier is working. Other areas are also getting a lot better, so all of this improvement is very unlikely to be from the air purifier.

  B111A North small table B111A Outside clean rooms west of north small table B111B Control area B111B Outside clean room by northeast shelves B111B Inside clean room B111 Antechamber
0.3 um 0 46 88 59 0 1182
0.5 um 0 9 42 28 0 283
1.0 um 0 4 23 20

0

165

The particle counter in the B111B clean room read 14/0/0 on a first reading, so I more carefully turned it on by staying under the particle counter and it then read 0/0/0.

 

We installed the 2.75" 1550 nm AR Coated Viewport onto the Laser Filter Cavity (LFC) Output Vacuum Cube with an annealed copper gasket. The gasket had some discoloration/very light scratches on one face, but I think this is ok. We tightened the screws loosely, then tightly by hand, then by hand using the socket on the screws, then using the wrench with my hand near the bolt, then using the full length of the wrench, then using the torque wrench at 6 Nm, 7 Nm, 8 Nm, and finally at 9.6 Nm. 9.6 Nm was enough so that the flange appeared to be fully tightened. Special care was used to ensure we didn't touch the glass of the viewport flange. 14 Nm was used for the 2.75" blank flange and regular copper gasket.

We then pushed the LFC Output Vacuum Cube into place, secured it to the optics tables, and brought the LFC Input Vacuum Cube closer to the edge of the table so that we can more easily work on it. We tightened that cube to the table as well. 

GQuEST Output Filter Cavities
Torrey Cullen - posted 15:16, Tuesday 14 January 2025 (12075)
First Cavity Lock post 2025 Eaton Canyon Fire

We had been discussing if the extremely poor air quality due to the fire had effected our superoptics in the OFCs. To check this I performed a couple quick ringdown measurements on OFC1. For both wavelengths, the finesse of the cavity is close to measurements before the fires. It seems uneffected. For reference F = pi * c / L * (measured tau in screenshot).

Images attached to this report
Environmental Dust/Particulates
Daniel Grass - posted 14:00, Tuesday 14 January 2025 (12074)
Particle Count Measurements on January 14, 2025 in B111, B111A, B111B

I took some particle counts today with the full 1 minute capture time. They are much better than yesterday.

B111A North large table North small table South large table South small table Outside clean rooms in middle of room south of curtain Outside clean rooms west of north small table before air purifier was turned on  Outside clean rooms west of north small table after air purifier was turned on for ~7 minutes
0.3 um 9 1 411 1 56 145 81
0.5 um 2 0 50 0 19 52 26
1.0 um 0 0 13 0 12

34

17

See attached photo for the location of the air purifier (with a newly installed filter and run at fan setting 3 with UV on) and the particle counter for the last two measurements.

 

B111B Control area Outside clean room by northeast shelves Inside clean room
0.3 um 3526 77 1
0.5 um 1598 48 0
1.0 um 579 26 0

The 0.3 um particle happened within the first second of the minute capture time, so might be been caused by me setting it up (I didn't have a beard net or gloves on).

 

Other B111 antechamber
0.3 um 1516
0.5 um 489
1.0 um 235

 

For each respective location outside the clean rooms, these are some of the lowest particle counts we have taken. Particle counts are a bit elevated inside the clean rooms.

Images attached to this report
Lasers Amplifiers
Torrey Cullen - posted 11:33, Tuesday 14 January 2025 (12073)
Lasers on in B111B

Laser and amplifier are being turned on following 12065. After the work done on the interlock system it does not seem like the signs are functional yet. 

Environmental Dust/Particulates
Daniel Grass - posted 14:43, Monday 13 January 2025 (12071)
Particle Count Measurements on January 13, 2025 in B110, B111, B111A, B111B, and B111D

I took some particle counts today with the full 1 minute capture time 

B111A North large table North small table South large table South small table Outside clean rooms in middle of room Outside clean rooms west of north small table
0.3 um 6893 9768 5309 3362 15614 15383
0.5 um 873 1098 595 436 2075 2088
1.0 um 339 395 214 156 877 854

Something is clearly wrong inside the clean rooms. Also quite elevated (~20x) from September outside clean room. Please note the exposure time was 1/6th as long in the September measurments.

 

B111B Control area Outside clean room by northeast shelves Inside clean room
0.3 um 18352 4460 0
0.5 um 2433 655 0
1.0 um 971 305 0

~2x elevated from September outside the clean room by the northeast shelves.

 

Other B111 antechamber B111D
0.3 um 17033 13006
0.5 um 3581 2214
1.0 um 1965 1104

~10x elevated from September in B111 antechamber. First measurements in B111D to my knowledge.

 

B110 Outside next to fume hood Inside fume hood
0.3 um 14887 244330
0.5 um 2555 2441
1.0 um 1251 1169

Yes, the 0.3 um particle count was that high. First measurements in B110 to my knowledge. 

Environmental Dust/Particulates
Torrey Cullen - posted 18:12, Sunday 12 January 2025 - last comment - 19:57, Sunday 12 January 2025(12067)
Particle Readings after January 2025 Eaton Canyon Fire

[Torrey,Jeff,Ian]

We decided to check on air quality in bridge/lab. The hallways and offices are suprisingly decent. As you enter the lab it is more and more smokey. Inside the cleanroom in B111B is by far the worst smelling. May need to replace the filters in the cleanroom fans. For now though, cleanroom1.jpg and cleanroom2.jpg show no particles above .3 um in the cleanroom.  Outside the cleanroom the particle counts seem higher than normal, would have to find old log posts with these numbers though.

Additionally we did a quick sweep in B111A. It smells like an animal has died in the vents somewhere. I will look into putting in a maintenence request tomorrow morning. Particle count where the smell is the strongest: deadanimal.jpg. Additionally the B111A cleanroom fans do not seem to be on. insideB111Acleanroom.jpg (this is on the table with the RQB stuff).

Images attached to this report
Comments related to this report
Daniel Grass - 19:43, Sunday 12 January 2025 (12068)

From September when we were moving in: https://mccullerlab.com/logs/lab/index.php?callRep=11873

Daniel Grass - 19:57, Sunday 12 January 2025 (12069)

These results are ~5x worse in B111B outside the clean room and ~30x worse in B111A outside the clean room. There could be some variance in room location, but surely not 30x unless the particle counter was in a really weird location or there was a lot of actively right before and during the count.

Non-image files attached to this comment
Environmental Dust/Particulates
Daniel Grass - posted 19:28, Tuesday 07 January 2025 (12066)
Smoke smell in B111

It smells like smoke in B111, probably from the Eaton Fire. I couldn't take a particle count since the particle counter was dead and I couldn't find the cord to charge it.

Vacuum Systems General (GQuEST)
Daniel Grass - posted 17:49, Monday 16 December 2024 - last comment - 16:41, Tuesday 14 January 2025(12054)
CF0600S2, 4.5 in blank flange, and 2.75 in blank flange added to LFC Vacuum Cube

[Jeff, Daniel]

I tightened the 10" to 4.5" zero length reducer flange on top of the vacuum cube.

We added the CF0600S2 6" to two 2.75" CF Multiport Reducer to the 10" to 6" CF Zero Length Reducer. We did the usual method of placing the flange using some top screws to hold it, adding 2 bottom screws, remove the top screws, tilting the flange up, dropping in the gasket, and then adding the remaining screws. Only a few screws were accessible with our torque wrenches, so I tightened the screws by hand in a star pattern (going every 5 screws with 16 total). We bought this part (as part of a set) so that the torque wrenches can reach all the screws.

We took off a plastic cover and cleaned the 10" to 4.5" reducer flange, which was covered with duct tape (and duct tape residue) outside where the 4.5" flange goes (and outside the vacuum). 

We added a 4.5" CF blank flange to this 10" to 4.5" reducer flange on top of the vacuum cube. Because there were only 8 screws, I started at a lower torque, 6 N m, did a star pattern every 3 screws so that every screw was tightened twice, and incremented the torque by 1 N m until the torque was 12 N m, then I incremented the torque by 3 N m until 34 N m when the flange seemed tight enough (stainless steel contact around ~70% of the surface).

We added a 2.75" CF blank flange to the CF0600S2 (where the laser view port will not go) and attached it with the plate nuts. I lightly torqued by hand and used the torque wrench starting at its minimum value of 6 N m. I tightened every 2-3 bolts until the torque wrench did not tighten the bolts further, then incremented the torque by 1 N m until 14 N m. For the laser viewport, we will use the annealed copper gasket, which will require less torque to tighten.

Next, we will swap out the 10" to 2.75" reducer flange for a 10" to 4.5" reducer flange because the 10" to 2.75" flange is needed for the demonstrator IFO.

Comments related to this report
Daniel Grass - 16:41, Tuesday 14 January 2025 (12077)

I forgot to note at the time that the gasket for the 6" flange had some discoloration/very light scratches on one face, but I think this is fine.

Environmental Dust/Particulates
Daniel Grass - posted 18:15, Thursday 19 September 2024 - last comment - 14:48, Monday 13 January 2025(11873)
Particle Count Measurements in B102, B111, B111A, and B111B

I took some particle counter measurments in B102, B111, B111A, and B111B. In all the clean rooms, the particle count was 0. These clean rooms were the B102 cleanroom with the optical cavities, the B111B mobile cleanrooms, the Southern most B111A long table, and the Northern most B111B short table.

I also took some 10 second particle counts outside the clean rooms. Data is 0.3 um count/0.5 um count/1.0 um count)

B102: 1268/72/28 (there were lots of people and activity during this measurment)

B111 (entryway): 495/46/28

B111B: 177/51/39

B111A: 95/17/13

These seem like very good measurments.

The particle counter is on the optics tables in B111B.

Comments related to this report
Daniel Grass - 19:58, Sunday 12 January 2025 (12070)

It should probably read "Northern most B111A short table" considering there is no short table in B111B.

Daniel Grass - 14:48, Monday 13 January 2025 (12072)

Data extrapolated to a full minute for easy comparisons:

B102: 7608/432/168 (there were lots of people and activity during this measurment)

B111 (entryway): 2970/276/168

B111B: 1062/306/234

B111A: 564/102/78

Displaying reports 221-240 of 1079.Go to page Start 8 9 10 11 12 13 14 15 16 End